A person has $ 200 to spend on two goods X and Y whose respective prices are $10 and $20
a. Draw the budget line.
b. What happens to the original budget line if the budget falls by 25%?
c. What happens to the original budget line if the price of X doubles?
d. What happens to the original budget line if the price of Y falls to $4?
1
Expert's answer
2020-10-05T13:27:10-0400

a)Px=10

Gx= 200/10= 20

Py=20

Gy= 200/20=10

The budget line is as below.






b) If the budget falls by 25%.

200-(0.25*200)=150

Px=10,Gx=150/10 =15

Py=20,Gy=150/20=7.5

The new budget line is as shown below.






c)If the price of X doubles

Px=10*2=20,Gx=200/20=10

Py=20,Gy=200/20=10

The budget line is as below;





d)When the price of Y falls to $4.

Px=10,Gx=200/10=20

Py=4, Gy=200/4=50.

The budget line is as below;